# Compositional Difference-in-Differences for categorical outcomes

### v2

@@ 1.1.1. Analytical framework
A useful way to think about CTT is through a neutral individual who starts indifferent, assigning equal probability to all $p$ categories, $(1/p)$. After treatment, this person updates their weights according to CTT. Components greater than $1/q$ indicate categories that gain relative importance, while smaller components indicate categories that lose. For example, in a three-party voting scenario, a neutral voter starts with $(1/3,1/3,1/3)$. After a campaign (treatment), CTT might imply $(0.5,0.3,0.2)$. This shows that support for the first party increased, while the other two declined, with the second party still stronger than the third.
1.1.2. Parallel growths
@@
This section develops the identification strategy and clarifies its implications for the probability mass function. I start with the following assumption.
Assumption 1 (Common Support).
The support, denoted by $\bar{\mathcal{Y}}=\{c_{1},\dots,c_{p}\}$, is the same for all untreated potential categorical outcomes.
This assumption ensures that all observed and counterfactual outcomes in the absence of treatment are defined over a common set of categories. It rules out situations where the support changes across groups or over time, which would make comparisons and counterfactual reasoning impossible. In practice, it implies that the set of possible outcomes is fixed. Such support conditions are standard in the literature. To introduce the main identifying assumption, I provide intuition for the DiD identification logic. The core idea is that, in the absence of treatment, the treated group would have followed the same trajectory as the control group. Figure 2 illustrates this DiD identification logic: the control group, unexposed to the treatment, serves as a counterfactual benchmark, mirroring what would have happened to the treated group had it not received the intervention.
$q^{N}_{0,0}$$q^{N}_{1,0}$$q^{N}_{0,1}$$q^{N}_{1,1}$$q^{I}_{1,1}$ControlGroupTreatmentGroupTime 0Time 1TimeTrendcounterfactualpathActualtreatmentpathCausal effectTime Periods
@@
Figure 2. Visual representation of the DiD identification strategy. The counterfactual outcome for the treated group in the post-treatment period ($q^{N}_{1,1}$) is constructed by extrapolating the time trend observed in the control group (blue arrows) to the treated group’s pre-treatment level. The causal effect is the difference between the observed treated outcome ($q^{I}_{1,1}$) and this counterfactual ($q^{N}_{1,1}$), shown by the orange double-headed arrow.
@@
As I said before, in many empirical settings, the sizes of the groups defined in each category can vary dramatically in scale. Therefore, a standard parallel trends assumption on the raw quantities is problematic in this context, as it would impose the same absolute change on both large and small groups, which is often unrealistic. To address this issue of scale and to formulate a more plausible identifying assumption, I instead consider parallel trends on the log transformation of the quantities. This approach focuses on proportional, rather than absolute, changes. For a given vector, define the functions $(\log\cdot)$ and $(\exp\cdot)$ which apply respectively the logarithm and the exponential functions component-wise to the vectors.
@@
$\log\cdot\left(q_{1,1}^{N}\right)-\log\cdot\left(q_{1,0}^{N}\right)=\log\cdot\left(q_{0,1}^{N}\right)-\log\cdot\left(q_{0,0}^{N}\right).$
Assumption 2 is the core identifying condition for the model. It states that, in the absence of treatment, the proportional growth (or decline) of each category in the treated group would have been the same as that of the control group. This reflects the intuitive idea that underlying trends affect all categories proportionally to their size. This may appear as a more realistic and flexible assumption than one based on raw quantities in many economic contexts, especially when the sizes of categories differ significantly. It captures the idea that trends often affect all units proportionally to their size. It is also possible to use discrete covariates to make the assumption more believable. I discuss this possibility in appendix B. From this assumption, the counterfactual quantity vector for the treated group can be easily recovered, and the corresponding counterfactual probability mass can be obtained by normalization, as I see in Theorem 1. Taken together, assumptions 1 and 2 impose sufficient structure to point-identify the counterfactual distributions. The following theorem 1 formalizes this result
@@
$q_{1,1}^{N}=\exp\cdot\left(\log\cdot\left(q_{1,0}^{N}\right)+\log\cdot\left(q_{0,1}^{N}\right)-\log\cdot\left(q_{0,0}^{N}\right)\right),$ (3)
$S_{1,1}^{N}=\sum_{k=1}^{p}\exp\cdot\left(\log\cdot\left(q_{1,0}^{N}(c_{k})\right)+\log\cdot\left(q_{0,1}^{N}(c_{k})\right)-\log\cdot\left(q_{0,0}^{N}(c_{k})\right)\right),$ (4)
$\pi_{1,1}^{N}=\frac{1}{S_{1,1}^{N}}q_{1,1}^{N}.$ (5)
@@
This identification result is constructive and provides a straightforward way to implement the method. In the next section, I will discuss further justifications for parallel growths, exploring its economic and geometric implications.
1.2. Economic justifications of parallel growths
Consider a standard Random Utility Model, in which a decision-maker selects one alternative from a finite set of mutually exclusive choices $\{c_{1},c_{2},\dots,c_{p}\}$. The utility that the individual derives from alternative $c_{k}$ in the absence of treatment is modeled as:
$U^{N}_{gt}(c_{k})=V^{N}_{gt}(c_{k})+\varepsilon_{gt}(c_{k}),$
where $U^{N}_{gt}(c_{k})$ denotes the latent utility of category $k$ in group $g$ at time $t$ in the absence of treatment, and $V^{N}_{gt}(c_{k})$ is the systematic (deterministic) component of utility (can be seen as the measure of the total attractiveness of the category) and $\varepsilon_{gt}(c_{k})$ are the random unobserved shocks. The individual is assumed to choose the alternative that yields the highest latent utility:
$\text{Choice}=\arg\max_{c_{k}}U^{N}_{gt}(c_{k}).$
I decompose the total attractiveness into two parts:
$V^{N}_{gt}(c_{k})=\underbrace{\mu^{N}_{gt}(c_{k})-\log\left(\sum_{k}e^{\mu^{N}_{gt}(c_{k})}\right)}_{\text{relative attractiveness}}+\underbrace{\log\left(S^{N}_{gt}\right)}_{\text{absolute attractiveness}}.$
This decomposition shows that $V^{N}_{gt}(c_{k})$ contains two additive components: the first is the relative desirability of the category (within-choice structure) $\mu^{N}_{gt}(c_{k})-\log\Big(\sum_{j}e^{\mu_{gt}(c_{j})}\Big)$. This term reflects how desirable option $c_{k}$ is relative to all other alternatives at the same $(g,t)$. The second is absolute desirability (aggregate intensity) $\log(S_{gt})$. This term captures the overall size or participation level (e.g., how many people are making a choice at all). It scales all utilities equally, affecting the total mass of choices but not their relative shares. When it rises (say, due to higher purchasing power, stronger participation, or favorable aggregate conditions), all utilities shift up equally, people are more likely to choose something, but the proportions among categories might remain the same unless the first component also change. Together, they form a utility concept that blends relative preferences and absolute participation. This decomposition is useful to see how both components can be affected. We finally have:
$U^{N}_{gt}(c_{k})=\mu^{N}_{gt}(c_{k})-\log\left(\sum_{k}e^{\mu^{N}_{gt}(c_{k})}\right)+\log\left(S^{N}_{gt}\right)+\varepsilon_{gt}(c_{k}).$
In the traditional RUM, we assume that the random terms $(\varepsilon_{gt}(c_{1}),\dots,\varepsilon_{gt}(c_{p}))$ all follow standard type-1 extreme value distribution and are independent. Because each marginal error term remains Gumbel-distributed with mean $\gamma\approx 0.5772$ (the Euler–Mascheroni constant), the expected utility of alternative $c_{k}$ is:
$\mathbb{E}[U^{N}_{gt}(c_{k})]=\mu^{N}_{gt}(c_{k})-\log\left(\sum_{k}e^{\mu^{N}_{gt}(c_{k})}\right)+\log\left(S^{N}_{gt}\right)+\gamma.$
$\mathbb{E}[U^{N}_{gt}(c_{k})]$ captures the average latent total desirability in the population for category $c_{k}$. The first implication of the RUM is given by the following proposition.
Proposition 1 (Implication for expected utilities).
$\mathbb{E}[U^{N}_{11}(c_{k})]-\mathbb{E}[U^{N}_{10}(c_{k})]=\mathbb{E}[U^{N}_{01}(c_{k})]-\mathbb{E}[U^{N}_{00}(c_{k})]\quad\forall\,c_{k}\in\bar{\mathcal{Y}}.$ (6)
This equivalence provides a clear economic interpretation of the identifying assumption. It states that, in the absence of treatment, both groups would have experienced the same evolution in how desirable each option is on average. For example, in a voting context, suppose the control group experiences a rise in the expected utility of voting Democrat (e.g., due to national political trends). My assumption implies that the treated group would have experienced the same increase in expected utility had it not been exposed to the policy. The same logic applies to all categories. Thus, the parallel growths assumption is not merely a mathematical condition on outcomes; it can also be seen as an economic hypothesis about the common evolution of underlying category-specific attractiveness in the absence of treatment. This provides a micro-founded justification for CoDiD: it estimates how a policy shifts the total attractiveness of alternatives as measured by expected utility.
@@ 1.2. Economic justifications of parallel growths
Now, we will see how it also restrict the evolution of the shares. Define the multinomial logistic transform. For $\pi=[\pi(c_{1}),\dots,\pi(c_{p})]\in\mathcal{S}^{p-1}$. The log-odds (or multinomial logit) transformation $\ell:\mathcal{S}^{p-1}\to\mathbb{R}^{p-1}$ maps probabilities to real-valued indices:
$\ell(\pi)=\Bigg[\log\left(\frac{\pi(c_{1})}{\pi(c_{p})}\right),\dots,\log\left(\frac{\pi(c_{p-1})}{\pi(c_{p})}\right)\Bigg].$ (7)
@@
Each component of this vector is the log-odds of category $c_{k}$ compared to the baseline category $c_{p}$. In fact, the ratio $\pi_{g,t}^{N}(c_{k})/\pi_{g,t}^{N}(c_{p})$ is the odds of category $c_{k}$ relative to the reference category $c_{p}$ and taking the logarithm gives the log-odds. The choice of the baseline category can be arbitrary. Its inverse is given by the softmax function: for any $y=(y_{1},\dots,y_{p-1})\in\mathbb{R}^{p-1}$,
@@
$\ell(\pi_{1,1})-\ell(\pi_{1,0})=\ell(\pi_{0,1})-\ell(\pi_{0,0}).$
@@
Moreover, parallel trends in log-odds is equivalent to parallel trends in utility differences for any pair of categories under the RUM:
$\Big\{\mathbb{E}[U^{N}_{11}(c_{k})]-\mathbb{E}[U^{N}_{11}(c_{l})]\Big\}-\Big\{\mathbb{E}[U^{N}_{10}(c_{k})]-\mathbb{E}[U^{N}_{10}(c_{l})]\Big\}=$
$\Big\{\mathbb{E}[U^{N}_{01}(c_{k})]-\mathbb{E}[U^{N}_{01}(c_{l})]\Big\}-\Big\{\mathbb{E}[U^{N}_{00}(c_{k})]-\mathbb{E}[U^{N}_{00}(c_{l})]\Big\},\quad\forall\,c_{k},c_{l}\in\bar{\mathcal{Y}}.$ (9)
@@
The log-odds of choosing one category over another provide a direct empirical analogue to differences in the underlying expected utilities. They summarize how much more desirable one option is relative to another, capturing the relative strength or simply the preferences in the population. For instance, in a voting context, the log-odds of choosing Democrat over Republican quantify the population’s relative preference for Democrats. When the log-odds rise, it means that the expected utility, or average attractiveness, of voting Democrat has increased relative to that of voting Republican. The parallel growths assumption implies that, in the absence of treatment, these relative preferences would have evolved in parallel across groups for any pair of categories. That is, any secular change, such as a national political shift increasing the attractiveness of Democrats relative to Republicans, would have affected both treated and control groups identically had the treatment not occurred. The treatment effect is then identified as the deviation from this common trajectory. From that, I can also identify the counterfactual distribution with an alternative formula:
$\pi_{1,1}^{N}=\ell^{-1}\left(\ell(\pi_{1,0})+\ell(\pi_{0,1})-\ell(\pi_{0,0})\right).$ (10)
@@
In a $\times2$ DiD with categorical outcomes, I have shown that assuming parallel trends in log-quantity ratios is not ad hoc, it is equivalent to assuming parallel trends in expected utilities, a natural behavioral assumption under the Random Utility Model. Because utility differences (not levels) drive choice, this assumption is economically meaningful. This foundation justifies the CoDiD approach as a method for estimating how a policy changes the relative attractiveness of each alternative compared to others. In Appendix D, I extend this result to more general Random Utility Models that allow for dependence among the error terms, thereby relaxing the IIA assumption implicitly assumed in the standard multinomial RUM.
1.3. Geometric justifications of parallel growths

### v3

@@ 1.1. Analytical framework
A useful way to think about CTT is through a neutral individual who starts indifferent, assigning equal probability to all $p$ categories $(1/p)$. After treatment, this person updates their weights according to CTT. Components greater than $1/P$ indicate categories that gain relative importance, while smaller components indicate categories that lose relative importance. For example, in a three-party voting scenario, a neutral voter starts with $(1/3,1/3,1/3)$. After a campaign (treatment), CTT might imply $(0.5,0.3,0.2)$. This shows that support for the first party increased, while the other two declined, with the second party still stronger than the third.
One can also consider the treatment effects on some functionals of the distribution. Consider, for example, a policy evaluated using a scalar index that is a nonlinear functional of compositional shares, such as the Herfindahl–Hirschman Index (HHI):
$\text{HHI}=\sum_{k=1}^{K}\pi(c_{k})^{2}$
A common approach would be to apply DiD directly to the index. However, this strategy extrapolates the index directly, ignoring that it is derived from shares that must satisfy some mathematical and economic constraints. A better approach is to get the counterfactual shares first and later recompute the counterfactual index. Consider a general function $H$ on the shares, a causal analysis on $H(\pi)$ would be the functional treatment effect on the treated (FTT), defined for a given choice of $H$
$\text{FTT}=H(\pi_{1,1}^{1})-H(\pi_{1,1}^{0}).$ (5)
In the next paragraph, I introduce the identifying assumption and its implications.
1.2. Parallel growths
@@
This section develops the identification strategy and clarifies its implications for the shares. I start with the following assumption.
Assumption 1 (Positivity).
For all $g,t$, $q^{0}_{g,t}>0$
Assumption 1 guarantees that all category-specific quantities are strictly positive in every group and period in the absence of treatment. This ensures that both observed and counterfactual compositional vectors are well defined and comparable across groups and time. In particular, it rules out settings where a category disappears entirely for one group or in one period, which would break the alignment of supports required for the DiD analysis. Assumptions of this form are standard in compositional settings.
I now turn to the core identification idea. The logic mirrors the standard DiD framework: absent treatment, the treated group would have evolved in parallel with the control group. Figure 1 illustrates this intuition; the untreated control group traces out the counterfactual trajectory that the treated group would have followed had the intervention not occurred.
$q^{0}_{0,0}$$q^{0}_{1,0}$$q^{0}_{0,1}$$q^{0}_{1,1}$$q^{1}_{1,1}$ControlGroupTreatmentGroupTime 0Time 1TimeTrendcounterfactualpathActualtreatmentpathCausal effectTime Periods
@@
Figure 1. Visual representation of the DiD identification strategy. The counterfactual outcome for the treated group in the post-treatment period ($q^{0}_{1,1}$) is constructed by extrapolating the time trend observed in the control group (blue arrows) to the treated group’s pre-treatment level. The causal effect is the difference between the observed treated outcome ($q^{1}_{1,1}$) and this counterfactual ($q^{0}_{1,1}$), shown by the orange double-headed arrow.
@@
In many empirical settings, the sizes of the groups defined in each category can vary dramatically in scale. Therefore, a standard linear parallel trends assumption on the raw quantities can be problematic in this context, as it would impose the same absolute change on both large and small groups, which is often unrealistic. To address this issue of scale and to formulate a more plausible identifying assumption, I instead consider parallel trends on the log transformation of the quantities. This approach focuses on proportional, rather than absolute, changes. For a given vector, define the functions $(\log)$ and $(\exp)$ which apply respectively the logarithm and the exponential functions component-wise to the vectors.
@@
$\log\left(q_{1,1}^{0}\right)-\log\left(q_{1,0}^{0}\right)=\log\left(q_{0,1}^{0}\right)-\log\left(q_{0,0}^{0}\right).$
Assumption 2 is the core identifying condition for the model. It states that, in the absence of treatment, the proportional growth (or decline) of each category in the treated group would have been the same as that of the control group. This reflects the intuitive idea that underlying trends affect all categories proportionally to their size. This may appear as a more realistic and flexible assumption than one based on raw quantities in many economic contexts, especially when the sizes of categories differ significantly. It captures the idea that trends often affect all units proportionally to their size. It is also possible to use discrete covariates to make the assumption more believable. I discuss this possibility in appendix B. From this assumption, the counterfactual quantity vector for the treated group can be easily recovered, and the corresponding counterfactual probability mass can be obtained by normalization, as we can see in Theorem 1. Taken together, assumptions 1 and 2 impose sufficient structure to point-identify the counterfactual distributions. The following theorem 1 formalizes this result
@@
$q_{1,1}^{0}=\exp\left(\log\left(q_{1,0}^{0}\right)+\log\left(q_{0,1}^{0}\right)-\log\left(q_{0,0}^{0}\right)\right),$ (6)
$S_{1,1}^{0}=\sum_{k=1}^{p}\exp\left(\log\left(q_{k,1,0}^{0}\right)+\log\left(q_{k,0,1}^{0}\right)-\log\left(q_{k,0,0}^{0}\right)\right),$ (7)
$\pi_{1,1}^{0}=\frac{1}{S_{1,1}^{0}}q_{1,1}^{0}.$ (8)
@@
This identification result is constructive and provides a straightforward way to implement the method.Before I present the economic and geometric implication of parallel growths, I show how this assumption restricts the evolution of each margin.
1.2.1. Restrictions on margins
@@
The parallel growths assumption that we impose directly on categorical quantities also restrict the evolution of shares and totals in a certain way. To see how it restricts the evolution of the shares, define the multinomial logistic transform. $\pi=[\pi_{1},\dots,\pi_{p}]\in\mathcal{S}^{q-1}$. The log-odds (or multinomial logit) transformation $\ell:\mathcal{S}^{p-1}\to\mathbb{R}^{p-1}$ maps probabilities to real-valued indices:
$\ell(\pi)=\Bigg[\log\left(\frac{\pi_{1}}{\pi_{p}}\right),\dots,\log\left(\frac{\pi_{p-1}}{\pi_{p}}\right)\Bigg].$
@@
Each component of this vector is the log-odds of category $k$ compared to the baseline category $p$. In fact, the ratio $\pi_{k,g,t}^{0}/\pi_{p,g,t}^{0}$ is the odds of category $k$ relative to the reference category $p$, and taking the logarithm gives the log-odds. The choice of the baseline category can be arbitrary. Its inverse is given by the softmax function: for any $y=(y_{1},\dots,y_{p-1})\in\mathbb{R}^{p-1}$,
@@
$\ell(\pi^{0}_{1,1})-\ell(\pi^{0}_{1,0})=\ell(\pi^{0}_{0,1})-\ell(\pi^{0}_{0,0}).$
From that, I can also identify the counterfactual distribution with an alternative formula:
$\pi_{1,1}^{0}=\ell^{-1}\left(\ell(\pi_{1,0})+\ell(\pi_{0,1})-\ell(\pi_{0,0})\right).$ (10)
We are also interested in examining how it restricts the evolution of the totals. Re-expressing the parallel growths assumption in terms of observable totals and category shares yields a transparent decomposition:
$\log(S_{1,1}^{0})-\log(S_{1,0}^{0})=\underbrace{\log(S_{0,1}^{0})-\log(S_{0,0}^{0})}_{\text{Control-group total growth}}+\underbrace{\log\left[\sum_{k=1}^{p}\left(\frac{\pi_{k,0,1}^{0}}{\pi_{k,0,0}^{0}}\right)\pi_{k,1,0}^{0}\right]}_{\text{Composition adjustment factor }\lambda}.$ (11)
Here, $S_{1,0}^{0}$ is the treated group’s observed pre-treatment total, $S_{0,1}^{0}/S_{0,0}^{0}$ is the aggregate growth rate in the control group, and
$\lambda\equiv\sum_{k=1}^{p}\left(\frac{\pi_{k,0,1}^{0}}{\pi_{k,0,0}^{0}}\right)\pi_{k,1,0}^{0}$
is the composition adjustment factor, a reallocation multiplier that captures how the treated group’s initial allocation across categories aligns with the control group’s internal reallocation dynamics. The factor $\lambda$ is essential because it accounts for a fundamental source of heterogeneity that direct DiD on log-total ignores: even when two groups experience identical category-level growth rates, they will exhibit different aggregate outcomes if their initial compositions differ. Specifically: If the treated group is initially concentrated in categories that gained relative share in the control group ($\pi_{k,0,1}^{0}>\pi_{k,0,0}^{0}$), then $\lambda>1$, implying a counterfactual total larger than what aggregate scaling alone would suggest. If it is concentrated in categories that lost relative share, then $\lambda<1$, implying a counterfactual decline, even if the control group’s total is unchanged. Only when $\pi_{1,0}^{0}=\pi_{0,0}^{0}$ does $\lambda=1$, reducing Equation (11) to the standard DiD in log-total prediction.
To see why the composition adjustment factor matters, consider evaluating an early voting policy. Suppose the control state shows no change in total turnout across elections, but its vote composition shifts: Democratic vote share rises while Republican share falls. Now assume the treated state, before the policy, has an electorate that is disproportionately Republican, the group losing relative share in the control state. Even with flat aggregate turnout in the control state, the treated state’s counterfactual turnout would have declined absent the policy, because its composition is weighted toward the shrinking group. In CoDiD, this appears as a composition adjustment factor. Standard DiD, which imposes parallel trends in total turnout, would wrongly assume no counterfactual change and thus underestimate the policy’s effect.
1.2.2. Economic justifications of parallel growths
Consider a standard Random Utility Model, in which a decision-maker selects one alternative from a finite set of mutually exclusive choices $\{1,2,\dots,p\}$. The utility that the individual derives from alternative $k$ in the absence of treatment is modeled as:
$U^{0}_{k,g,t}=V^{0}_{k,g,t}+\varepsilon^{0}_{k,g,t},$
where $U^{0}_{k,g,t}$ denotes the latent utility of category $k$ in group $g$ at time $t$ in the absence of treatment, and $V^{0}_{k,g,t}$ is the systematic (deterministic) component of utility (can be seen as the measure of the total attractiveness of the category) and $\varepsilon^{0}_{k,g,t}$ are the random unobserved shocks. The individual is assumed to choose the alternative that yields the highest latent utility:
$\text{Choice}=\arg\max_{k}U^{0}_{k,g,t}.$
I decompose the total attractiveness into two parts:
$V^{0}_{k,g,t}=\underbrace{\mu^{0}_{k,g,t}}_{\text{relative attractiveness}}-\underbrace{\log\left(\sum_{k}e^{\mu^{0}_{k,g,t}}\right)+\log\left(S^{0}_{g,t}\right)}_{\text{absolute attractiveness}}.$
This decomposition shows that $V^{0}_{k,g,t}$ contains two additive components: the first is the relative desirability of the category (within-choice structure) $\mu^{0}_{k,g,t}$. This term reflects how desirable option $k$ is relative to all other alternatives at the same $(g,t)$. The second is absolute desirability (aggregate intensity) $\log\Big(\sum_{j}e^{\mu^{0}_{j,g,t}}\Big)+\log(S^{0}_{g,t})$. This term captures the overall size or participation level (e.g., how many people are making a choice at all). It scales all utilities equally, affecting the total mass of choices but not their relative shares. When it rises (say, due to higher purchasing power, stronger participation, or favorable aggregate conditions), all utilities shift up equally, people are more likely to choose something, but the proportions among categories might remain the same unless the first component also changes. Together, they form a utility concept that blends relative preferences and absolute participation. This decomposition is useful to see how both components can be affected. We finally have:
$U^{0}_{k,g,t}=\mu^{0}_{k,g,t}-\log\left(\sum_{k}e^{\mu^{0}_{k,g,t}}\right)+\log\left(S^{0}_{g,t}\right)+\varepsilon^{0}_{k,g,t}.$
In the traditional RUM, we assume that the random terms $(\varepsilon^{0}_{1,g,t},\dots,\varepsilon^{0}_{p,g,t}$ all follow standard type-1 extreme value distribution and are independent. Because each marginal error term remains Gumbel-distributed with mean $\gamma\approx 0.5772$ (the Euler–Mascheroni constant), the expected utility of alternative $k$ is:
$\mathbb{E}[U^{0}_{k,g,t}]=\mu^{0}_{k,g,t}-\log\left(\sum_{k}e^{\mu^{0}_{k,g,t}}\right)+\log\left(S^{0}_{g,t}\right)+\gamma.$
$\mathbb{E}[U^{0}_{k,g,t}]$ captures the average latent total desirability in the population for category $k$. The first implication of the RUM is given by the following proposition. Define the following vector of expected utilities:
$\mathbb{E}[U^{0}_{g,t}]=\Big[\mathbb{E}[U^{0}_{1,1,1}],\cdots,\mathbb{E}[U^{0}_{p,1,1}]\Big].$
Proposition 2 (Implication for expected utilities).
$\mathbb{E}[U^{0}_{1,1}]-\mathbb{E}[U^{0}_{10}]=\mathbb{E}[U^{0}_{01}]-\mathbb{E}[U^{0}_{00}].$ (12)
This equivalence provides a clear economic interpretation of the identifying assumption. It states that, in the absence of treatment, both groups would have experienced the same evolution in how valuable each option is on average. For example, in a voting context, suppose the control group experiences a rise in the expected utility of voting Democrat (e.g., due to national political trends). Assumption  2 means that the treated group would have experienced the same change in expected utility had it not been exposed to the policy. The same logic applies to all categories. Thus, the parallel growths assumption is not merely a mathematical condition on outcomes; it can also be seen as an economic hypothesis about the common evolution of underlying category-specific attractiveness in the absence of treatment. This provides a micro-founded justification for CoDiD: it estimates how a policy shifts the attractiveness of alternatives as measured by expected utility.
Proposition 3 (Implication for shares).
@@ 1.2.2. Economic justifications of parallel growths
Parallel trends in log-odds is equivalent to parallel trends in utility differences for any pair $k,j$ of categories under the RUM:
$\Big\{\mathbb{E}[U^{0}_{k,1,1}]-\mathbb{E}[U^{0}_{j,1,1}]\Big\}-\Big\{\mathbb{E}[U^{0}_{k,1,0}]-\mathbb{E}[U^{0}_{j,1,0}]\Big\}=$
$\Big\{\mathbb{E}[U^{0}_{k,0,1}]-\mathbb{E}[U^{0}_{j,0,1}]\Big\}-\Big\{\mathbb{E}[U^{0}_{k,0,0}]-\mathbb{E}[U^{0}_{j,0,0}]\Big\}$ (13)
@@
The log-odds of choosing one category over another provide a direct empirical analogue to differences in the underlying expected utilities. They summarize how much more desirable one option is relative to another, capturing the relative strength or simply the preferences in the population. For instance, in a voting context, the log-odds of choosing Democrat over Republican quantify the population’s relative preference for Democrats. When the log-odds rise, it means that the expected utility, or average attractiveness, of voting Democrat has increased relative to that of voting Republican. The parallel growths assumption implies that, in the absence of treatment, these relative preferences would have evolved in parallel across groups for any pair of categories. That is, any secular change, such as a national political shift increasing the attractiveness of Democrats relative to Republicans, would have affected both treated and control groups identically had the treatment not occurred. The treatment effect is then identified as the deviation from this common trajectory.
@@
To summarize, I have shown that assuming parallel trends in log-quantity ratios is not ad hoc; it is equivalent to assuming parallel trends in expected utilities, a natural behavioral assumption under the Random Utility Model. Because utility differences (not levels) drive choice, this assumption is economically meaningful. This foundation justifies the CoDiD approach as a method for estimating how a policy changes the relative attractiveness of each alternative compared to others. In Appendix D, I extend this result to more general Random Utility Models that allow for dependence among the error terms, thereby relaxing the IIA assumption implicitly assumed in the standard multinomial RUM.
1.2.3. Geometric justifications of parallel growths